# Revisiting Randomization in Greedy Model Search

### v1

@@ 6 Explaining the Role of Randomization in RFs
Taken together, these results broaden our perspective on the role of randomization in the success of RFs. Whereas Mentch and Zhou [42] concluded that feature subsampling is only useful in low SNR regimes, these recent results show that it can also be useful even in the absence of any noise [18]. Furthermore, since feature subsampling is popularly viewed as a form of implicit regularization, these results urge us to pay more attention to nuance in our understanding of this phenomenon, which has become of great interest to the machine learning community [53, 43, 3, 4]. Specifically, our mental models of how regularization works in ridge and lasso regression may not map neatly onto the effects of algorithmic regularization, be it from feature subsampling, dropout, or gradient descent.
7 Discussion and Future Work
The major theoretical limitations of our work are our assumptions that the features are orthogonal. This gap is somewhat bridged by our empirical results, which show that RGS continues to provide performance gains even in correlated settings, such as under banded and block covariance structures. Nonetheless, a promising direction for future work is to explain these results theoretically. For example, one could study how small deviations from orthogonality (e.g., through structured perturbations of $\mathbf{X}^{\top}\mathbf{X}/n$) affect the RGS weights and associated bias-variance tradeoffs. Second, while we make arguments about random forests, we are only able to do so by analogy, having analyzed feature subsampling in RGS. A direct theoretical analysis of how feature subsampling improves RFs is therefore of significant interest. Partial results were obtained by Mei et al. [40], but they require certain oracle assumptions, which make their analysis deviate from how RFs are used in practice.
In addition, RGS faces certain practical limitations. While RGS significantly reduces computational complexity via our dynamic programming implementation, its efficiency gains diminish as $k$ grows large relative to $p$. Investigating further computational optimizations—such as parallelized implementations or approximate search heuristics—would be beneficial for scaling to ultra-high-dimensional datasets.
By addressing these challenges, RGS has the potential to become a broadly useful tool for sparse modeling, particularly in high-stakes domains that demand both accuracy and interpretability. Promising application areas include high-dimensional causal inference and deep learning pipelines as a feature selection module.
Acknowledgments
@@
Appendix A Proofs
A.1 Proofs for Finite RGS Weights
@@
    If $i\neq j$, the remaining $k-1$ steps are equivalent to running RGS on $p-1$ features. The coefficients are reassigned as follows:
@@
      If $i<j$, which occurs with probability $\mathbb{P}(J<j)$, the index $j$ is shifted to the $(j-1)$-th position in the new ordering. The weight contribution becomes $w_{j-1}^{k-1,m,p-1}$.
At the first step of RGS, we select the feature with the largest absolute inner product $|\hat{\beta}_{j,p}|=|\langle\mathbf{y},\mathbf{x}_{\cdot j,p}\rangle|$ from an $m$-sized subset of $p$ features. Because there are no ties in the feature rankings, this is equivalent to selecting the minimum index $j$ in this subset. The probability of $x_{j,p}$ being selected is thus the probability that $j$ is the minimum index in a random $m$-sized subset of $p$ elements, which is $\binom{p-j}{m-1}/\binom{p}{m}$. Therefore, by the law of total probability,
$w_{j}^{k,m,p}$ $=\mathbb{P}(J=j)+w_{j}^{k-1,m,p-1}\mathbb{P}(J>j)+w_{j-1}^{k-1,m,p-1}\mathbb{P}(J<j)$ (21)
@@
Now, by the hockey-stick identity from combinatorics, we can simplify the coefficients from (21):
@@
Plugging in these two expressions into (21) gives us recurrence (12).
To prove monotonicty of the weights with respect to $j$, recall that $w_{j}^{k,m,p}=\mathbb{P}(x_{j,p}\in\mathcal{M}_{k})$. Here, the randomness is over the choices of the sets, $\boldsymbol{{\mathcal{V}}}\coloneqq\{\mathcal{V}_{1},\mathcal{V}_{2},\ldots,\mathcal{V}_{k}\}$. Define the events
$\mathcal{E}_{1}\coloneqq\{x_{j,p}\in\mathcal{M}_{k},\;x_{j-1,p}\notin\mathcal{M}_{k}\},$
and
$\mathcal{E}_{2}\coloneqq\{x_{j-1,p}\in\mathcal{M}_{k},\;x_{j,p}\notin\mathcal{M}_{k}\}.$
To define $T$, consider $\boldsymbol{\mathcal{V}}\in\mathcal{E}_{1}$. For this random seed, there is some index $1\leq i\leq k$ such that $x_{j,p}$ is selected at the $i$-th iteration. This implies that we have $x_{j-1,p}\notin\mathcal{V}_{i}$, otherwise, $x_{j-1,p}$ would have been selected in the $i$-th step instead. Furthermore, we have $x_{j,p}\notin\mathcal{V}_{l}$ for $l>i$. Define $T(\boldsymbol{\mathcal{V}})\coloneqq\{\mathcal{V}_{1},\ldots,\mathcal{V}_{i-1},\mathcal{V}_{i}^{\prime},\mathcal{V}_{i+1}^{\prime},\ldots,\mathcal{V}_{k}^{\prime}\}$, where $\mathcal{V}_{i}^{\prime}=\mathcal{V}_{i}\cup\{x_{j-1,p}\}\backslash\{x_{j,p}\}$, and for $l>i$, $\mathcal{V}_{l}^{\prime}=\mathcal{V}_{l}$ if $x_{j-1,p}\notin\mathcal{V}_{l}$ otherwise $\mathcal{V}_{l}^{\prime}=\mathcal{V}_{i}\cup\{x_{j,p}\}\backslash\{x_{j-1,p}\}$. Denote $\boldsymbol{\mathcal{V}}^{\prime}\coloneqq T(\boldsymbol{\mathcal{V}})$ for convenience. Let $x_{j_{l},p}$ and $x_{j^{\prime}_{l},p}$ be the features selected at the $l$-th step with the set of subdictionaries $\boldsymbol{\mathcal{V}}$ and $\boldsymbol{\mathcal{V}}^{\prime}$, respectively. Note that $x_{j_{l},p}=x_{j^{\prime}_{l},p}$ for $l=1,\ldots,i-1$, since the algorithm only makes use of $\mathcal{V}_{1},\ldots,\mathcal{V}_{i-1}$ to select these terms. Similarly, $x_{j_{l},p}=x_{j^{\prime}_{l},p}$ for $l=i+1,\ldots,k$, because the combinatorics depend solely on the ordinal structure of the coefficients $|\hat{\beta}_{l,p}|$. Thus, $x_{j-1,p}$ in $\boldsymbol{\mathcal{V}}^{\prime}$ plays the same role as $x_{j,p}$ in $\boldsymbol{\mathcal{V}}$. For the same reason, $\mathbb{P}(\boldsymbol{\mathcal{V}})=\mathbb{P}(\boldsymbol{\mathcal{V}}^{\prime})$. We also see that $\boldsymbol{\mathcal{V}}^{\prime}\in\mathcal{E}_{2}$, which proves that $T$ is well-defined and is measure-preserving. ∎
A.2 Proofs for Asymptotic RGS Weights
@@
$\tilde{w}_{j}^{k,\gamma}=(e^{\alpha}-1)e^{-\alpha j}+e^{-\alpha j}\tilde{w}_{j}^{k-1,\gamma}+\big{(}1-e^{-\alpha(j-1)}\big{)}\tilde{w}_{j-1}^{k-1,\gamma}.$ (22)
However, recurrence (22) contains dependencies on both $\tilde{w}_{j}^{k-1,\gamma}$ and $\tilde{w}_{j-1}^{k-1,\gamma}$. These dependencies make it difficult to construct a closed form and uncover properties of $\tilde{w}_{j}^{k,\gamma}$. Ideally, we would like recurrences involving a change in $j$ to be separate from recurrences involving a change in $k$. Fortunately, recurrence (22) can be separated into the following pair of recurrences.
Lemma A.2.
From the the recurrence in Lemma A.1, we can construct the following recurrence pair for $j\geq 1$:
$\tilde{w}_{j}^{k,\gamma}=\big{(}e^{-\alpha(j-k)}-e^{-\alpha j}\big{)}\big{(}1-\tilde{w}_{j}^{k-1,\gamma}\big{)},\quad\tilde{w}_{j}^{0,\gamma}=0,$ (23)
$\tilde{w}_{j}^{k,\gamma}=1-e^{-\alpha k}-\big{(}e^{-\alpha(k-j+1)}-e^{-\alpha k}\big{)}\tilde{w}_{j-1}^{k,\gamma},\quad\tilde{w}_{0}^{k,\gamma}=0.$ (24)
Proof of Lemma A.1.
To show the limit form of the coefficients in recurrence (22), we derive the following result using basic asymptotic facts. For a fixed $j$, for $m,p\to\infty$ and $m/p\to\gamma$, we have
@@
Since $\tilde{w}_{j}^{k,\gamma}$ and $\tilde{w}_{j-1}^{k,\gamma}$ exist for all $j$ by inductive assumption, $w_{j}^{k+1,\gamma}$ exists as well for all $j$. Therefore, by induction, we know that the limit $\tilde{w}_{j}^{k,\gamma}$ exists for all $j,k$ and satisfies the recurrence in Lemma A.1. ∎
Proof of Lemma A.2.
@@
For recurrence (23), we plug in $\tilde{w}_{j}^{0,\gamma}=0$ on the right hand side:
@@
Thus, the recurrence (23) satisfies the base case $k=1$. Now, note that $\tilde{w}_{j-1}^{1,\gamma}=(e^{\alpha}-1)e^{-\alpha(j-1)}$. Plugging in this expression into recurrence (24), we see that
@@
Consider step $(k+1)$. We show that $\tilde{w}_{j}^{k+1,\gamma}$ satisfies recurrence (23) in the following way. We start with the recurrence from Lemma A.1:
@@
Substituting $\tilde{w}_{j-1}^{k,\gamma}$ using recurrence (24) under the inductive assumption, we obtain the following result:
@@
Therefore, recurrence (23) is satisfied for step $(k+1)$. Now, we prove that recurrence (24) is satisfied for step $(k+1)$ in a similar fashion. We plug in recurrence (23) into Lemma A.1:
@@
Thus, recurrence (24) is satisfied for step $(k+1)$. Therefore, by induction, for a fixed $j$, both recurrences are satisfied for all $k\in\mathbb{N}^{+}$. ∎
@@
To prove property (a), note that from Lemma A.1, we have that $\tilde{w}_{j}^{k,\gamma}$ exists. From (23), a first-order non-homogeneous recurrence relation with feature coefficients, we directly obtain the following closed form for $\tilde{w}_{j}^{k,\gamma}$ [24, Equation 1.2.4]:
@@
To prove property (b), we first show the lower bound. According to Lemma A.5, we have $\tilde{w}_{j}^{k,\gamma}\geq\tilde{w}_{j}^{k-1,\gamma}$. From (23), we know
@@
Alternatively, we can rewrite recurrence (23) as
@@
$\begin{split}\frac{1-e^{-\alpha k}}{1-e^{-\alpha k}+e^{-\alpha(k-j)}}&=\frac{1}{1+\left(1-e^{-\alpha k}\right)^{-1}e^{-\alpha(k-j)}}\\ &=\frac{1}{1+\left(e^{\alpha k}-1\right)^{-1}e^{aj}}\\ &=\frac{1}{1+e^{-\alpha(h(a,k)-j)}}.\qed\end{split}$ (25)
@@
To prove property (ii), note that from (22) we have
@@
Finally, we re-index by setting $i$ to be $k-i$, yielding
@@
We start to prove Theorem 4.2. For the following proofs, we perform the substitution $d=k-j$.
@@
$\check{w}_{d}^{\gamma}=\sum^{\infty}_{i=1}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}.$ (26)
Proof of Lemma A.4.
@@
Note that by substituting (23) into (24), we have
@@
Thus, $\tilde{w}_{j+1}^{k,\gamma}-\tilde{w}_{j}^{k-1,\gamma}=e^{-\alpha k}\big{(}\tilde{w}_{j}^{k,\gamma}-\tilde{w}_{j}^{k-1,\gamma}\big{)}\geq 0$ for all $j,k\geq 1$, where we use Lemma A.5. Fix $d$, let $j=k-d$, and consider the sequence $\big{\{}\tilde{w}_{k-d}^{k,\gamma}\big{\}}^{\infty}_{k=d+1}$. This sequence is bounded and non-decreasing; thus, by the monotone convergence theorem, the limit of $\big{\{}\tilde{w}_{k-d}^{k,\gamma}\big{\}}^{\infty}_{k=d+1}$ exists. Hence, $\check{w}_{d}^{\gamma}$ exists and we must now identify its limit.
We will prove the closed form (26) by showing that for every $\epsilon>0$, there exists some $K$ such that for all $k>K$,
@@
First, fix $\epsilon$ and suppose $k$ is even. We begin by using the triangle inequality, substituting in the closed form for $\tilde{w}_{k-d}^{k,\gamma}$ we obtained in Lemma A.3:
@@
$\leq\sum^{k}_{i=k/2+1}e^{\alpha i(d-i/2+1/2)}\left(1-\prod^{k}_{l=k-i+1}\big{(}1-e^{-l\alpha}\big{)}\right)+\sum^{k/2}_{i=1}e^{\alpha i(d-i/2+1/2)}\left(1-\prod^{k}_{l=k-i+1}\big{(}1-e^{-l\alpha}\big{)}\right).$ (27)
Suppose that $k>2d$. We bound the first sum of (27). By basic inequalities, we have
@@
When $k\to\infty$, this final expression will go 00, so there exists some $K_{1}$ such that for all $k>K_{1}$, the final expression is less than $\epsilon/2$.
We perform a similar computation to bound the second sum in expression (27). Note that
@@
$1-\prod^{k}_{l=k-i+1}\big{(}1-e^{-l\alpha}\big{)}\leq 1-(1-e^{-(k/2+1)\alpha})^{k/2}\leq\frac{k}{2}\cdot e^{-(k/2+1)\alpha}.$ (28)
Substituting expression (28) back into the second sum of expression (27), we have that
@@
Note that as $k\to\infty$, this final expression goes to 00, so there exists some $K_{2}$ such that for all $k>K_{2}$, this expression is less than $\epsilon/2$.
@@
$\sum^{2d}_{i=1}(-1)^{i+1}e^{-(\alpha/2)(i-(d+1/2))^{2}}=\sum^{d}_{i=1}(-1)^{i+1}\big{(}e^{-(\alpha/2)(i-(d+1/2))^{2}}-e^{-(\alpha/2)((d+1/2)-i)^{2}}\big{)}=0.$ (29)
@@
$\sum^{2d+1}_{i=0}(-1)^{i+1}e^{-(\alpha/2)(i-(d+1/2))^{2}}=0.$ (30)
Therefore, using the cancellation of (29) and (30),

### v3

@@ 8 Discussion
These results, alongside recent related work, broaden our understanding of randomization in RFs. mentch2020randomization concluded that feature subsampling is mainly useful in low-SNR regimes, but our results show it can also help in the absence of noise (see Proposition 6.1). Moreover, since feature subsampling is often viewed as implicit regularization, our findings urge closer attention to the nuances of this phenomenon. In particular, mental models derived from ridge and lasso regression may not transfer cleanly to algorithmic regularization via feature subsampling, dropout, or gradient descent (wager2013dropout; neyshabur2015searchrealinductivebias; arora2019implicit; bartlett2020benign).
The main theoretical limitation of our work is the orthogonality assumption. Our simulations partially bridge this gap, showing that EFS continues to improve upon FS in non-orthogonal settings. Extending the theory to handle correlated features is a promising direction for future work. For instance, one could study how structured perturbations of $\mathbf{X}^{\top}\mathbf{X}/n$ affect the ensemble weights. Additionally, our analysis applies directly only to forward selection. Extending the theoretical framework of mei2024exogenous to characterize how feature subsampling improves actual RFs remains an important open problem.
Reproducibility
The code used to reproduce all numerical experiments in this paper is available at
 github.com/yuyumi/EFS.
Acknowledgements
@@
Supplement to “Revisiting Randomization in Greedy Model Search”
S.1 Proofs
S.1.1 Proofs for finite EFS($k,m$) weights
Proof of Lemma 3.1.
To prove monotonicty of the weights with respect to $j$, recall that $w_{j}^{k,m,p}=\mathbb{P}(x_{j,p}\in\mathcal{M}_{k})$. Here, the randomness is over the choices of the sets, $\boldsymbol{{\mathcal{V}}}=\{\mathcal{V}_{1},\mathcal{V}_{2},\ldots,\mathcal{V}_{k}\}$. Define the events
$\mathcal{E}_{1}=\{x_{j,p}\in\mathcal{M}_{k},\;x_{j-1,p}\notin\mathcal{M}_{k}\},$
and
$\mathcal{E}_{2}=\{x_{j-1,p}\in\mathcal{M}_{k},\;x_{j,p}\notin\mathcal{M}_{k}\}.$
[…]
@@
    If $i\neq j$, the remaining k1k-1 steps are equivalent to running EFS on p1p-1 features. The coefficients are reassigned as follows:
@@
      If $i<j$, which occurs with probability $\mathbb{P}(J<j)$, the index $j$ is shifted to the (j1)(j-1)-th position in the new ordering. The weight contribution becomes $w_{j-1}^{k-1,m,p-1}$.
At the first step of EFS, we select the feature with the largest absolute inner product $|\hat{\beta}_{j,p}|=|\langle\mathbf{y},\mathbf{x}_{\cdot j,p}\rangle|$ from an $m$-sized subset of $p$ features. Because there are no ties in the feature rankings, this is equivalent to selecting the minimum index $j$ in this subset. The probability of $x_{j,p}$ being selected is thus the probability that $j$ is the minimum index in a random $m$-sized subset of $p$ elements, which is $\binom{p-j}{m-1}/\binom{p}{m}$. Therefore, by the law of total probability,
$w_{j}^{k,m,p}$ $=\mathbb{P}(J=j)+w_{j}^{k-1,m,p-1}\mathbb{P}(J>j)+w_{j-1}^{k-1,m,p-1}\mathbb{P}(J<j)$ (S.1)
@@
Now, by the hockey-stick identity from combinatorics, we can simplify the coefficients from (S.1):
@@
Plugging in these two expressions into (S.1) gives us recurrence (8). ∎
S.1.2 Proofs for asymptotic EFS($k,m$) weights
@@
$\tilde{w}_{j}^{k,\gamma}=(e^{\alpha}-1)e^{-\alpha j}+e^{-\alpha j}\tilde{w}_{j}^{k-1,\gamma}+\big(1-e^{-\alpha(j-1)}\big)\tilde{w}_{j-1}^{k-1,\gamma}.$ (S.2)
However, recurrence (S.2) contains dependencies on both $\tilde{w}_{j}^{k-1,\gamma}$ and $\tilde{w}_{j-1}^{k-1,\gamma}$. These dependencies make it difficult to construct a closed form and uncover properties of $\tilde{w}_{j}^{k,\gamma}$. Ideally, we would like recurrences involving a change in $j$ to be separate from recurrences involving a change in $k$. Fortunately, recurrence (S.2) can be separated into the following pair of recurrences.
Lemma S.1.2.
From the the recurrence in Lemma S.1.1, we can construct the following recurrence pair for $j\geq 1$:
$\tilde{w}_{j}^{k,\gamma}=\big(e^{-\alpha(j-k)}-e^{-\alpha j}\big)\big(1-\tilde{w}_{j}^{k-1,\gamma}\big),\quad\tilde{w}_{j}^{0,\gamma}=0,$ (S.3)
$\tilde{w}_{j}^{k,\gamma}=1-e^{-\alpha k}-\big(e^{-\alpha(k-j+1)}-e^{-\alpha k}\big)\tilde{w}_{j-1}^{k,\gamma},\quad\tilde{w}_{0}^{k,\gamma}=0.$ (S.4)
Proof of Lemma S.1.1.
To show the limit form of the coefficients in recurrence (S.2), we derive the following result using basic asymptotic facts. For a fixed $j$, for $m,p\to\infty$ and $m/p\to\gamma$, we have
@@
Since $\tilde{w}_{j}^{k,\gamma}$ and $\tilde{w}_{j-1}^{k,\gamma}$ exist for all $j$ by inductive assumption, $w_{j}^{k+1,\gamma}$ exists as well for all $j$. Therefore, by induction, we know that the limit $\tilde{w}_{j}^{k,\gamma}$ exists for all $j,k$ and satisfies the recurrence in Lemma S.1.1. ∎
Proof of Lemma S.1.2.
@@
For recurrence (S.3), we plug in $\tilde{w}_{j}^{0,\gamma}=0$ on the right hand side:
@@
Thus, the recurrence (S.3) satisfies the base case $k=1$. Now, note that $\tilde{w}_{j-1}^{1,\gamma}=(e^{\alpha}-1)e^{-\alpha(j-1)}$. Plugging in this expression into recurrence (S.4), we see that
@@
Consider step $(k+1)$. We show that $\tilde{w}_{j}^{k+1,\gamma}$ satisfies recurrence (S.3) in the following way. We start with the recurrence from Lemma S.1.1:
@@
Substituting $\tilde{w}_{j-1}^{k,\gamma}$ using recurrence (S.4) under the inductive assumption, we obtain the following result:
@@
Therefore, recurrence (S.3) is satisfied for step $(k+1)$. Now, we prove that recurrence (S.4) is satisfied for step $(k+1)$ in a similar fashion. We plug in recurrence (S.3) into Lemma S.1.1:
@@
Thus, recurrence (S.4) is satisfied for step $(k+1)$. Therefore, by induction, for a fixed $j$, both recurrences are satisfied for all $k\in\mathbb{N}^{+}$. ∎
@@
To prove property (a), note that from Lemma S.1.1, we have that $\tilde{w}_{j}^{k,\gamma}$ exists. From (S.3), a first-order non-homogeneous recurrence relation with feature coefficients, we directly obtain the following closed form for $\tilde{w}_{j}^{k,\gamma}$ (elaydi2005introduction, Equation 1.2.4):
@@
To prove property (b), we first show the lower bound. According to Lemma S.1.5, we have $\tilde{w}_{j}^{k,\gamma}\geq\tilde{w}_{j}^{k-1,\gamma}$. From (S.3), we know
@@
Alternatively, we can rewrite recurrence (S.3) as
@@
$\begin{split}\frac{1-e^{-\alpha k}}{1-e^{-\alpha k}+e^{-\alpha(k-j)}}&=\frac{1}{1+\left(1-e^{-\alpha k}\right)^{-1}e^{-\alpha(k-j)}}\\ &=\frac{1}{1+\left(e^{\alpha k}-1\right)^{-1}e^{aj}}\\ &=\frac{1}{1+e^{-\alpha(h(a,k)-j)}}.\qed\end{split}$ (S.5)
@@
To prove property (ii), note that from (S.2) we have
@@
Finally, we re-index by setting $i$ to be kik-i, yielding
@@
We start to prove Theorem 3.2. For the following proofs, we perform the substitution d=kjd=k-j.
@@
$\check{w}_{d}^{\gamma}=\sum^{\infty}_{i=1}(-1)^{i+1}e^{\alpha i(d-i/2+1/2)}.$ (S.6)
Proof of Lemma S.1.4.
@@
Note that by substituting (S.3) into (S.4), we have
@@
Thus, $\tilde{w}_{j+1}^{k,\gamma}-\tilde{w}_{j}^{k-1,\gamma}=e^{-\alpha k}\big(\tilde{w}_{j}^{k,\gamma}-\tilde{w}_{j}^{k-1,\gamma}\big)\geq 0$ for all $j,k\geq 1$, where we use Lemma S.1.5. Fix $d$, let j=kdj=k-d, and consider the sequence $\big\{\tilde{w}_{k-d}^{k,\gamma}\big\}^{\infty}_{k=d+1}$. This sequence is bounded and non-decreasing; thus, by the monotone convergence theorem, the limit of $\big\{\tilde{w}_{k-d}^{k,\gamma}\big\}^{\infty}_{k=d+1}$ exists. Hence, $\check{w}_{d}^{\gamma}$ exists and we must now identify its limit.
We will prove the closed form (S.6) by showing that for every $\epsilon>0$, there exists some $K$ such that for all $k>K$,
@@
First, fix $\epsilon$ and suppose $k$ is even. We begin by using the triangle inequality, substituting in the closed form for $\tilde{w}_{k-d}^{k,\gamma}$ we obtained in Lemma S.1.3:
@@
$\leq\sum^{k}_{i=k/2+1}e^{\alpha i(d-i/2+1/2)}\left(1-\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big)\right)+\sum^{k/2}_{i=1}e^{\alpha i(d-i/2+1/2)}\left(1-\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big)\right).$ (S.7)
Suppose that $k>2d$. We bound the first sum of (S.7). By basic inequalities, we have
@@
When $k\to\infty$, this final expression will go 0, so there exists some $K_{1}$ such that for all $k>K_{1}$, the final expression is less than $\epsilon/2$.
We perform a similar computation to bound the second sum in expression (S.7). Note that
@@
$1-\prod^{k}_{l=k-i+1}\big(1-e^{-l\alpha}\big)\leq 1-(1-e^{-(k/2+1)\alpha})^{k/2}\leq\frac{k}{2}\cdot e^{-(k/2+1)\alpha}.$ (S.8)
Substituting expression (S.8) back into the second sum of expression (S.7), we have that
@@
Note that as $k\to\infty$, this final expression goes to 0, so there exists some $K_{2}$ such that for all $k>K_{2}$, this expression is less than $\epsilon/2$.
@@
$\sum^{2d}_{i=1}(-1)^{i+1}e^{-(\alpha/2)(i-(d+1/2))^{2}}=\sum^{d}_{i=1}(-1)^{i+1}\big(e^{-(\alpha/2)(i-(d+1/2))^{2}}-e^{-(\alpha/2)((d+1/2)-i)^{2}}\big)=0.$ (S.9)
@@
$\sum^{2d+1}_{i=0}(-1)^{i+1}e^{-(\alpha/2)(i-(d+1/2))^{2}}=0.$ (S.10)
Therefore, using the cancellation of (S.9) and (S.10),